OpenStudy (anonymous):

Use graphical methods to solve the linear programming problem. 40x + 80y ≤ 560. 6x + 8y ≤ 72. x ≥ 0. y ≥ 0. Maximize P = 8x + 12y A. Maximum of 92 when x = 4 and y = 5 B. Maximum of 96 when x = 9 and y = 2 C. Maximum of 100 when x = 8 and y = 3 D. Maximum of 120 when x = 3 and y = 8

OpenStudy (anonymous):

I would substitute each and every option and see if it satisfies. ex. P= 8(3)+12(8)= 96+24=120 (check) but it fails to satisfy 40x+80y=560, etc...

OpenStudy (anonymous):

If you want to solve this graphically, you need to draw two axis, X and Y. In these axis you display the four equations that you've listed, where you've replaced the inequality with equality, so 40x+80y=560. For each line that you've drawn, you have to determine which part of the graph suits your constraint. Mark that part with a color or so. Your solution is the area which has four colors, meaning: the area which is valid for all of the four equations. See the drawing for the graphical approach.

OpenStudy (anonymous):